# Quantum memory assisted observable estimation

### v3

@@ Appendix A CPS method
where $M_{q}$ is the amount of repreparations of $|{\Phi_{N}}\rangle$. Finally, the variance (43) can be written as follows:
$\sigma^{2}\hat{{\Phi_{a}}}\approx\frac{3+\cos{(8{\Phi}_{N})}}{16M_{q}}\leq\frac{1}{4M_{q}}.$ (46)
From (38) one can conclude that
$\sin{\Phi_{N}}-\tan{\Phi_{a}}\cos{\Phi_{N}}=R(\tan{\Phi_{a}}\cos{\Theta_{R}}-\sin{\Theta_{R}})$ (47)
The linear combination of sine and cosine waves is equivalent to a single sine wave with a phase shift and scaled amplitude
$-\sqrt{1+\tan^{2}{\Phi_{a}}}\cos{\left(\Phi_{N}+\arctan{\left(\frac{1}{\tan{\Phi_{a}}}\right)}\right)}=R(\tan{\Phi_{a}}\cos{\Theta_{R}}-\sin{\Theta_{R}}).$ (48)
Then the target angle is
$\Phi_{N}$ $=$ $\arccos{\left(\frac{R\cos{\Theta_{R}}(\tan{\Theta_{R}}-\tan{\Phi_{a}})}{\sqrt{\tan^{2}{\Phi_{a}}+1}}\right)}-\arctan{\left(\frac{1}{\tan{\Phi_{a}}}\right)}.$
Assuming that $\Theta_{R}<<\Phi_{a}$, we can write
$\Phi_{N}$ $\sim$ $\arccos{\left(-\frac{N\varepsilon_{QSP}\tan{\Phi_{a}}}{\sqrt{\tan^{2}{\Phi_{a}}+1}}\right)}-\arctan{\left(\frac{1}{\tan{\Phi_{a}}}\right)}.$
Similarly to (43), one can deduce that the variance of the estimate of ${\Phi}_{N}$ scales as
$\sigma^{2}\hat{{\Phi}}_{N}$ $\sim$ $\frac{\left(N\epsilon_{QSP}+\sqrt{\frac{2-N^{2}\epsilon_{QSP}^{2}(\cos{2\Phi_{a}}-1)}{1+\cos{2\Phi_{a}}}}\right)^{2}(3+\cos{(8\Phi_{a})})\cot^{2}{\Phi_{a}}}{16M_{q}(1-N^{2}\epsilon_{QSP}^{2}+\cot^{2}{\Phi_{a}})}\sim\frac{1}{M_{q}},$ (49)
where we used that $|R|<N\varepsilon_{QSP}$, $\varepsilon_{QSP}<<1$.
We demand the standard deviation $\sqrt{\eta}$ of the estimate $\hat{O}$ to be greater than the QSP errors. We can conclude that
$\varepsilon_{QSP}\leq\frac{\sqrt{\eta}}{N},\quad n_{QSP}=O\left(\frac{\log{(\frac{N}{\sqrt{\eta}})}}{\log{\log{(\frac{N}{\sqrt{\eta}})}}}\right),$ (50)
hold.
Appendix B Phase Wrapping Control
The state encoded in the memory qubit contains the phase ${\Phi}_{N}$, accumulated by the $N$ rounds of encoding. However, the phase in our method is estimated only up to modulus $2\pi$. With the growth of $N$ the phase ${\Phi}_{N}\notin[-\pi,\pi)$ and we have no information about how many times $L$ the phase wraps around $2\pi$. One can see that a strategy of directly estimating the phase will fail unless it is already known to sit within a window of width $2\pi/k$, where k is significantly large. To overcome this problem, we can use the method introduced in [10] and improved in [15] for the purpose of gate calibration. We select $\epsilon={\epsilon}_{0}$ and introduce the notation
$\xi_{0}={\epsilon}_{0}P_{sum},\quad P_{sum}\equiv\sum\limits_{j=1}^{N}\text{sign}{(a_{j}\langle P_{j}\rangle)}|a_{j}\langle P_{j}\rangle|,$ (51)
such that $\xi_{0}<2\pi$, holds. Here we state the result:
Algorithm B.1
Given a target precision $\eta>0$, and numbers $\alpha,\gamma\in Z^{+}$, the algorithm outputting an estimate $\hat{P}_{sum}$ of ${P}_{sum}$ proceeds as follows:
  • Fix $d_{L}=[\log_{2}{1/\eta}]$.
  • For $l=1,2,3,\dots,d_{L}$:
    • Obtain an estimate $(\hat{\xi}_{l})$ for $\xi_{l}=2^{l}{\epsilon}_{0}P_{sum}\mod{2\pi}$ from $M_{l}=\alpha+\gamma(d_{L}-l)$ repetitions of of measurement circuit.
    • If $l=1$, set $\widehat{\epsilon_{0}P^{(1)}_{sum}}\equiv(\hat{\xi}_{0})$.
    • Else, set $\widehat{\epsilon_{0}{P}^{(l)}_{sum}}$ to be the (unique) number in $[\widehat{\epsilon_{0}{P}^{(l-1)}_{sum}}-\frac{\pi}{2^{l}},\widehat{\epsilon_{0}{P}^{(l-1)}_{sum}}+\frac{\pi}{2^{l}})$ such that
      $2^{l}\widehat{\epsilon_{0}{P}^{(l)}_{sum}}\equiv(\hat{\xi}_{l-1}),$ (52)
      holds.
  • Return $\hat{P}_{sum}=\hat{P}^{(d_{L})}_{sum}$ as an estimate for ${P}_{sum}$.
The error probability on every step is
$p\left(2^{l}\widehat{\epsilon_{0}{P}^{(l)}_{sum}}\right)=P\left(2^{l}(\widehat{\epsilon_{0}{P}^{(l)}_{sum}}-\epsilon_{0}{P}_{sum})\geq\frac{\pi}{2}\bigvee 2^{l}(\widehat{\epsilon_{0}{P}^{(l)}_{sum}}-\epsilon_{0}{P}_{sum})<-\frac{\pi}{2}\right).$ (53)
An error on every step will will lead to an incorrect range of the estimate. Following [15], this probability is bounded by $p\left(2^{l}\widehat{\epsilon_{0}{P}^{(l)}_{sum}}\right)\leq p_{max}\left(M_{l}\right)\sim M_{l}^{-1/2}2^{-M_{l}}$ and the variance of the estimate is
$\sigma^{2}\left(\widehat{\epsilon_{0}{P}^{(l)}_{sum}}\right)\leq(1-p_{max}\left(M_{d_{l}}\right))\frac{1}{2^{2(d_{l}+1)}}+\sum\limits_{l=1}^{d_{l}}p_{max}(M_{l})\frac{1}{2^{2l}}.$ (54)
The first term is the probability of no errors appear multiplied on the difference of the estimate from the true value on the final step and the second term is the probability of error appearing on every step multiplied on their contribution. Then the variance is bounded as
@@ Appendix B Phase Wrapping Control
$\sigma^{2}\left(\widehat{\epsilon_{0}{P}^{(l)}_{sum}}\right)\leq\frac{2^{\gamma-\alpha-2d_{l}}-2^{\gamma-\alpha-\gamma d_{l}}}{\left(2^{\gamma}-4\right)\sqrt{\gamma}}+\left(1-\frac{2^{-\gamma}}{\sqrt{\gamma}}\right)2^{-2(d_{l}+1)}.$ (55)
Choosing $\gamma>2$ and $\alpha>0$, prevents the sum from growing faster than $2^{-2d_{l}}$. The total amount of repetitions of the encoding process is denoted by $M_{q}=\sum\limits_{l=0}^{d_{L}}M_{l}$. Then
$\sigma^{2}\left(\widehat{\epsilon_{0}{P}^{(l)}_{sum}}\right)M_{q}\leq\left(-\frac{2^{-\gamma-2d_{l}-3}}{\sqrt{\gamma}}+\frac{2^{\gamma-\alpha-2d_{l}}-2^{\gamma-\alpha-\gamma d_{l}}}{\left(2^{\gamma}-4\right)\sqrt{\gamma}}+1\right)(2\alpha+\gamma(d_{l}-1))d_{l},$ (56)
and we can conclude that the variance on the final estimate scales like $\sim c(\gamma,\alpha)M_{q}^{-1}$, where $c$ is a constant. Choosing the constants $\alpha$ and $\gamma$ one can obtain better constant $c$. The detailed analyses one can find in [15].
The variance of the estimated observable by CPS method can be written as follows
$\sigma^{2}{(\hat{O}_{CPS})}\sim\frac{c(\gamma,\alpha)}{2^{2d_{L}}{\epsilon}_{0}^{2}M_{q}}.$ (57)
The total amount of state preparations for the CPS method is $T=Nn_{QSP}M_{q}+M_{1}=N(n_{QSP}M_{q}+n_{1})$. Here for every Pauli string we do $n_{QSP}$ QSP steps and we repeat this process for $N$ Pauli strings $M_{q}$ times. Finally, we do projective measurements to estimate the sign of every Pauli string $M_{1}$ times. Here $n_{1}$ is the amount of projective measurements per Pauli. The variance of the estimated observable scales as
$\sigma^{2}{(\hat{O}_{CPS})}\sim\frac{Cn_{QSP}}{\frac{T}{N}-n_{1}}\sim\frac{Cn_{QSP}N}{T},$ (58)
@@
$C=\frac{1}{{\epsilon_{0}^{2}}}\left(1-\frac{2^{-\gamma-4d_{l}-3}}{\sqrt{\gamma}}+\frac{2^{\gamma-\alpha-4d_{l}}-2^{\gamma-\alpha-\gamma d_{l}-2d_{L}}}{\left(2^{\gamma}-4\right)\sqrt{\gamma}}\right)(2\alpha+\gamma(d_{l}-1))d_{l}.$ (59)
We can select $\epsilon_{0}\sim 2^{-d_{l}}$ to hold. Then since (16) holds, we can see that $n_{QSP}\geq 2\max{|a_{j}|}$ By comparing the variances of the estimated observable by both methods, we get
$\frac{\sigma^{2}{(\hat{O}_{CPS})}}{\sigma^{2}{(\hat{O}_{QEE})}}\sim\frac{n_{QSP}}{N}.$ (60)
@@
It is possible to include additive errors into analyses of the probability (53). For example, the state preparation and measurement errors are of this type. Then performing two families of experiments, the probabilities of measuring $X=0$, $Y=0$ are
$P(X=0|\epsilon_{0}{P}_{sum},2^{l})=\frac{1}{2}(1+\cos{(2^{l}\epsilon_{0}{P}_{sum})})+\delta_{x}(2^{l}),\quad P(Y=0|\epsilon_{0}{P}_{sum},2^{l})=\frac{1}{2}(1-\sin{(2^{l}\epsilon_{0}{P}_{sum})})+\delta_{y}(2^{l}),$ (61)
@@
Appendix C Cost of implementing CPS

### v4

@@ Appendix A CPS method
where $M_{q}$ is the amount of repreparations of the state $|{\Phi_{N}}\rangle$. Finally, the variance (36) can be written as follows:
$\sigma^{2}(\hat{\Phi}_{N})\approx\frac{3+\cos{(8{\Phi}_{N})}}{16M_{q}}\leq\frac{1}{4M_{q}}.$ (39)
Note that this this variance is calculated for a fixed $\epsilon$ parameter. Then the variance of the estimate of the target observable is
$\sigma^{2}(\langle\hat{O}\rangle)=\sigma^{2}(\hat{\Phi}_{N})\epsilon^{-2}\sim\frac{1}{\epsilon^{2}M_{q}}.$ (40)
One can see that estimating the observable in this way we only achieve the standard quantum limit. Moreover, there is a problem of phase wrapping since we can only resolve phases mod $2\pi$. To resolve these issues, we adopt the technique of robust phase estimation from Refs. [13, 18] as detailed below.
Appendix B Phase Wrapping Control and Heisenberg Scaling
The state encoded in the memory qubit contains the phase
$\Phi=2\pi C+B,\quad C\in\mathrm{N},\quad B\in\mathrm{R},$ (41)
accumulated over $N$ rounds of encoding. In our method, since $\langle O\rangle$ is directly encoded as the rotation angle of a single qubit, expectation values differing in $C$ would be equivalent. An initial attempt to resolve this difficulty is to artificially scale $\langle O\rangle$ during the QSP step and rescale the estimate at the end of the protocol, using the phase estimation methods proposed by Higgins et al. [13] and later generalized by Kimmel et al. [18].
We introduce the notation
$\Phi_{l}(\epsilon_{0})\equiv\sum_{j=1}^{N}\tau^{(l)}_{j}(\epsilon_{0})\cos{(\theta_{j})},$ (42)
where we denote $\tau^{(l)}_{j}(\epsilon_{0})=2^{l-1}{\epsilon}_{0}s_{j}|a_{j}|$ such that $\Phi_{0}({\epsilon}_{0})\leq 2\pi$ holds. In  [18] the estimates $\hat{\Phi}_{l}(\epsilon_{0})$, $l=[1,d]$ of (42) are obtained from $M_{l}=\alpha+\gamma(d-l)$, $\gamma>2$, $\alpha>0$ repetitions of the measurement circuits described in the previous section. The probability of making an error on every step is
$p_{e}(\Phi_{l}(\epsilon_{0}))\equiv P\Big{[}\hat{\Phi}_{l}(\epsilon_{0})-\Phi_{l}(\epsilon_{0})\geq\frac{\pi}{2}\vee\hat{\Phi}_{l}(\epsilon_{0})-\Phi_{l}(\epsilon_{0})<-\frac{\pi}{2}\Big{]},$ (43)
and it is upper bounded by:
$p_{e}(\Phi_{l}(\epsilon_{0}))<\frac{1}{\sqrt{2\pi M_{l}}2^{M_{l}}}.$ (44)
On the last step $d=[\log_{2}{1/\sqrt{\eta}}]$ one obtains $\widehat{\epsilon_{0}\langle{O}\rangle}$ as an estimate of $\epsilon_{0}\langle O\rangle$ with a target variance $\eta$. For the total amount of state preparations $T$ the variance is $\sigma^{2}(\widehat{\epsilon_{0}\langle{O}\rangle})=\eta\sim O(T^{-2})$.
In contrast to the original method, where the initial encoded phase was assumed to be less than $2\pi$, we introduce a parameter $\epsilon_{0}$ that is sufficiently small to confine the phase within the desired range. However, this modification prevents direct application of the original algorithm, as the variance of the estimate $\langle\hat{O}\rangle$ would then scale as $O(\epsilon_{0}^{-2}T^{-2})=O(N^{2}/T^{2})$, preventing the attainment of Heisenberg scaling. To overcome this difficulty, we implement the described method in two stages, estimating $C$ and $B$ in (41) separately as it is shown below. Importantly, since QSP techniques allow us to encode each Pauli string as a phases between $\{0,2\pi\}$ with a nearly constant number of QSP iterations, we can estimate the observable $\hat{O}$ up to $O(1)$ variance using only $\log(N)$ state preparations for each Pauli string.
As we discussed in the introduction, we demand the overall QSP approximation error smaller than the target variance variance $\eta$ (i.e., $\varepsilon_{QSP}\leq\sqrt{\eta}/N$), so the number of steps $n_{QSP}$ in each QSP protocol is approximately constant with $N$ (see Eq. (25)). However, encoding each $\Phi_{l}({\epsilon}_{0})$ requires a different number of QSP steps to achieve a target error of $\varepsilon_{QSP}$. From condition (16), we derive the following constraint:
$n^{(l)}_{QSP}/2\geq|\tau_{l}(\epsilon_{0})|\equiv 2^{l-1}\epsilon_{0}\max\limits_{j=[1,N]}{|a_{j}|},\quad l\in\mathrm{N}.$ (45)
This is the minimal amount of the QSP steps needed to guarantee (22) and (23) to hold. Thus, for each $\Phi_{l}({\epsilon}_{0})$, the number of required QSP steps scales with $N$ as follows:
$n^{(l)}_{QSP}=O\left(|\tau_{l}(\epsilon_{0})|+\frac{\log{(\frac{N}{\sqrt{\eta}})}}{\log{\log{(\frac{N}{\sqrt{\eta}})}}}\right).$ (46)
In the first step of our procedure, we aim to estimate $C$, which represents the number of phase wraps in (41), with a target variance $\eta_{c}$. We assume that we wish to obtain an estimate of $\langle O\rangle$ with a variance $\eta\ll 1$ which means that $\eta_{c}\ll 1$. To estimate $C$, we perform $L=[\log_{2}{1/\sqrt{{\eta_{c}}}}]$ steps of the algorithm above and estimate $C$ as:
$\hat{C}=\Bigg{\lfloor}\frac{\hat{\Phi}_{L}(\epsilon_{0})}{2\pi}\Bigg{\rfloor}.$ (47)
If no errors occur in the principal range estimates for all $\Phi_{l}(\epsilon_{0})$, $l\in[1,L]$, then the maximum error in the estimate of $C$ is $2\pi/2^{L}$. We bound the variance of the estimate $\hat{C}$ of $C$ as:
$\sigma^{2}(\hat{C})\leq(1-p_{max}(M_{L}))\left(\frac{2\pi}{2^{L}}\right)^{2}+\sum\limits_{j=1}^{L}\left(\frac{2\pi}{2^{j}}\right)^{2}p_{max}(M_{j}).$ (48)
Doing the calculations, we get
$\sigma^{2}(\hat{C})\leq\frac{\pi^{2}}{4^{L}}\left(4+\frac{1}{\sqrt{2\pi\alpha}2^{\alpha}}\frac{16}{2^{\gamma}-2^{2}}\right).$ (49)
Note that the condition $\gamma>2$ prevents the sum growing faster then $4^{-L}$. The total number of estimates required for this stage is
$T_{c}=2N\sum\limits_{l=1}^{L}n^{(l)}_{QSP}M_{l}\leq 2Nn^{(1)}_{QSP}\sum\limits_{l=1}^{L}2^{l}M_{j}\leq 2^{L+1}Nn^{(1)}_{QSP}(\alpha+\gamma).$ (50)
Finally, the variance of the estimate $\hat{C}$ is
$\sigma^{2}(\hat{C})\leq\frac{N^{2}(n^{(1)}_{QSP})^{2}}{T_{c}^{2}}4\pi^{2}(\alpha+\gamma)^{2}\left(4+\frac{1}{\sqrt{2\pi\alpha}2^{\alpha}}\frac{16}{2^{\gamma}-2^{2}}\right).$ (51)
This approach ensures that the variance of the estimate is independent of $\epsilon_{0}$.
Now, we can apply the above algorithm to estimate $B\in(-\pi,\pi)$ with the target variance of $\eta_{b}\ll 1$. We directly encode $\Phi_{l}(1)=2^{l}P_{sum}$ by choosing $\epsilon_{0}=1$. The amount of QSP steps $\tilde{n}^{(l)}_{QSP}$ is given by (46) where $\epsilon_{0}=1$ holds. We perform $K=[\log_{2}{1/\sqrt{{\eta_{b}}}}]$ steps of the algorithm, and the estimate of $B$ is given by:
$\hat{B}=\hat{\Phi}_{K}(1)\mod 2\pi.$ (52)
The difference to the previous step is that even if all the steps of the algorithms provide no error, our final estimate can still differ from $B$ by at most $2\pi/2^{K+1}$. Then we bound the variance of the estimate $\hat{B}$ of $B$ as:
$\sigma^{2}(\hat{B})\leq(1-p_{max}(M_{K}))\left(\frac{2\pi}{2^{K+1}}\right)^{2}+\sum\limits_{l=1}^{K}\left(\frac{2\pi}{2^{l}}\right)^{2}p_{max}(M_{l}).$ (53)
Similarly to the previous step, we get
$\sigma^{2}(\hat{B})\leq\frac{\pi^{2}}{4^{K}}\left(1+\frac{1}{\sqrt{2\pi\alpha}2^{\alpha}}\left(3+\frac{16}{2^{\gamma}-2^{2}}\right)\right).$ (54)
The total amount of estimates required for this stage is
$T_{b}=2N\sum\limits_{l=1}^{K}\tilde{n}^{(l)}_{QSP}M_{l}\leq 2N\tilde{n}^{(1)}_{QSP}\sum\limits_{l=1}^{K}2^{l}M_{l}\leq 2^{K+1}N\tilde{n}^{(1)}_{QSP}(\alpha+\gamma),$ (55)
and the variance scales as
@@ Appendix B Phase Wrapping Control and Heisenberg Scaling
$\sigma^{2}(\hat{B})\leq\frac{N^{2}(\tilde{n}^{(1)}_{QSP})^{2}}{T_{b}^{2}}4\pi^{2}(\alpha+\gamma)^{2}\left(1+\frac{1}{\sqrt{2\pi\alpha}2^{\alpha}}\left(3+\frac{16}{2^{\gamma}-2^{2}}\right)\right).$ (56)
The total amount of state preparations needed for the CPS method is given by:
$T$ $=$ $Nn_{1}+T_{c}+T_{b}.$ (57)
Here, $Nn_{1}$ represents the small number of projective measurements performed beforehand to roughly estimate the mean values of each Pauli string, where $Nn_{1}<<T_{c},T_{b}$. As we write in the main text, this procedure is needed for the Pauli sign estimation. Then, using (50) and (55), we can write
$T\approx 2N(\alpha+\gamma)(2^{L}{n}^{(1)}_{QSP}+2^{K}\tilde{n}^{(1)}_{QSP}).$ (58)
The scaling of the QSP trials is the following
${n}^{(1)}_{QSP}=O\left(|\tau_{1}(\epsilon_{0})|+\frac{\log{(\frac{N}{\sqrt{\eta}})}}{\log{\log{(\frac{N}{\sqrt{\eta}})}}}\right)=O\left(\frac{1}{N}+\frac{\log{(\frac{N}{\sqrt{\eta}})}}{\log{\log{(\frac{N}{\sqrt{\eta}})}}}\right),$ (59)
$\tilde{n}^{(1)}_{QSP}=O\left(|\tau_{1}(1)|+\frac{\log{(\frac{N}{\sqrt{\eta}})}}{\log{\log{(\frac{N}{\sqrt{\eta}})}}}\right)=O\left(1+\frac{\log{(\frac{N}{\sqrt{\eta}})}}{\log{\log{(\frac{N}{\sqrt{\eta}})}}}\right).$
Then substituting $L$ and $K$, we get
@@
$T\approx 2N(\alpha+\gamma)\left(\frac{1}{\sqrt{\eta_{C}}}{n}^{(1)}_{QSP}+\frac{1}{\sqrt{\eta_{B}}}\tilde{n}^{(1)}_{QSP}\right).$ (60)
Since $C$ and $B$ are estimated independently, we can write
$\sigma^{2}(\hat{\Phi})=4\pi^{2}\sigma^{2}(\hat{C})+\sigma^{2}(\hat{B}).$ (61)
The target variance is $\eta$, so we can require equal contribution of variances: $\eta_{c}=\eta/(8\pi^{2})$, $\eta_{b}=\eta/2$. Then the variance of the target estimate is scaling as follows
$\eta\approx\frac{8N^{2}}{T^{2}}(2\pi{n}^{(1)}_{QSP}+\tilde{n}^{(1)}_{QSP})^{2}(\alpha+\gamma)^{2}.$ (62)
Hence, we can conclude that the variance of the estimate for $\langle O\rangle$ is
$\sigma^{2}(\hat{O}_{CPS})=\frac{8(\alpha+\gamma)^{2}N^{2}N^{2}_{QSP}}{T^{2}}=O\left(\frac{N^{2}N^{2}_{QSP}}{T^{2}}\right),$ (63)
where we used the notation $N_{QSP}\equiv 2\pi{n}^{(1)}_{QSP}+\tilde{n}^{(1)}_{QSP}=O\left(\frac{\log{(\frac{N}{\sqrt{\eta}})}}{\log{\log{(\frac{N}{\sqrt{\eta}})}}}\right)$.
@@
It is possible to include additive errors into analyses. For example, the state preparation and measurement errors are of this type. Then performing two families of experiments, the probabilities of measuring $X=0$, $Y=0$ are
$P(X=0|\epsilon_{0}{P}_{sum},2^{l})=\frac{1}{2}(1+\cos{(2^{l}\epsilon_{0}{P}_{sum})})+\delta_{x}(2^{l}),$ (64)
$P(Y=0|\epsilon_{0}{P}_{sum},2^{l})=\frac{1}{2}(1-\sin{(2^{l}\epsilon_{0}{P}_{sum})})+\delta_{y}(2^{l}),$
@@
Appendix C Comparison of CPS Method with Other Methods
To our best knowledge, the papers dedicated to the estimation of the mean of the complicated observable $O$ always use some variations of QEE or QPE methods [18, 32, 43]. That means, that every $\langle P_{i}\rangle$ is estimated individually, using one of these methods, and then summed classically. As we mentioned in the text, if each $\langle{P}_{j}\rangle$ is estimated with a variance $\sigma^{2}(\langle\hat{P}_{j}\rangle)=\eta_{p}$, the variance of the final estimate scales as $\sigma^{2}(\hat{O})\approx N\eta_{p}$, assuming roughly equal weights of the $\hat{P}_{j}$ in the decomposition of $\hat{O}$. That is why the only thing that can change from the method selection in the existing papers is the variance $\eta_{p}$.
In QEE every Pauli string is measured independently and then all the results are summed up. Hence, the variance of the estimate of $O$ is the following:
$\sigma^{2}{(\hat{O}_{QEE})}$ $\equiv\sum\limits_{i=1}^{N}\frac{a_{i}^{2}\sigma^{2}{\hat{P}_{i}}}{n_{c}}=\frac{N}{M_{c}}\sum\limits_{i=1}^{N}a_{i}^{2}\left(1-\langle\Psi_{0}|P_{i}|\Psi_{0}\rangle^{2}\right)\sim\frac{N^{2}}{M_{c}},$ (65)
where $M_{c}=Nn_{c}$ is the amount of projective measurements done in QEE. In QEE the amount of state preparations is equal to amount of measurements $T=M_{c}$. By comparing the variances of the estimated observable by both methods, we get
$\frac{\sigma^{2}{(\hat{O}_{CPS})}}{\sigma^{2}{(\hat{O}_{QEE})}}\sim\frac{N^{2}_{QSP}}{T}.$ (66)
The QPE scaling can significantly vary from the method. In [13, 18] the Heisenberg scaling is achieved, where the variance of one Pauli mean scales as $\sigma^{2}(\hat{P}_{QPE_{h}})\sim c^{2}/T_{P}^{2}$. Then the total variance is $\sigma^{2}(\hat{O}_{QPE_{h}})\sim Nc^{2}/T_{P}^{2}=N^{3}c^{2}/T^{2}$. Here $T=NT_{P}$ denotes the total amount of controlled-$U$ used to estimate one Pauli. CPS beats this method
$\frac{\sigma^{2}{(\hat{O}_{CPS})}}{\sigma^{2}{(\hat{O}_{QPE_{h}})}}\sim\frac{N^{2}_{QSP}}{N}.$ (67)
The Heisenberg QPE requires implementation of the observable as a unitary (in contrast to CPS) and enough circuit depth of the device. If one is able to implement the whole observable $O$, then the Heisenberg scaling for its estimate is achieved.
In [44] a special coefficient $\alpha\in[0,1]$ is introduced to weight both methods. When $\alpha=0$ the method is nothing else as QEE, when $\alpha=1$ the method is QPE. When $\alpha\in(0,1)$ some mean Pauli values are estimated according to QEE, some to QPE with $\alpha$ weights. One can see that if one have enough coherence time the variance of the method can’t be better then QPE’s, even if multiplied on some $\alpha$ dependent coefficient. Also, the result is again obtained for one Pauli, so in the end the total variance is formed by direct summation of $N$ variances for each Pauli estimate.
Knowing the standard deviation or the variance allows an immediate estimate of the accuracy of the estimated value. The Chebyshev’s inequality can be written
$\mathbb{P}\!\left(\Big{|}\hat{O}-{O}\Big{|}\geqslant\nu\!\right)\!\!\leqslant\frac{\sigma^{2}(\hat{O})}{\nu^{2}},$ (68)
where $\nu>0$. We use the Chebyshev’s inequality, since it does not require any assumptions on the probability distribution of the observable to be on a bounded support, but only existence of the finite variance. One can deduce that the amount of state preparations that guarantee the probability on the left hand side to be $p_{0}\in(0,1)$ if one uses different methods:
$T_{CPS}\sim\frac{NN_{QSP}}{\nu\sqrt{p_{0}}},\quad T_{QEE}\sim\frac{N^{2}}{\nu^{2}p_{0}},\quad T_{QPE_{h}}\sim\frac{N^{3/2}}{\nu\sqrt{p_{0}}}.$ (69)
The CPS outperforms all methods without paying a price of possibility to implement $O$ as a unitary that in general is not realistic for big $N$.
Appendix D Cost of implementing CPS